# A Survey of Neural Network Variational Monte Carlo from a Computing Workload Characterization Perspective

### 18 Mar 2026 v1

@@ I Introduction
NNVMC represents the electronic wavefunction with a neural network and optimizes model parameters under the variational principle. Prior studies show that this approach can achieve strong accuracy while offering favorable asymptotic scaling in representative settings (often cited as $O(N^{4})$ versus $O(N^{7})$ for CCSD(T)) [4]. Since Neural Quantum States were introduced, multiple neural wavefunction ansätze have been proposed. In this work, we focus on four representative models for ground-state problems: FermiNet [27], PauliNet [9], Psiformer [33], and Orbformer [6]. Other models, e.g., DeepSolid [20] and FiRE [30], are out of scope for this survey; in this study, we treat them as extensions of these parent families rather than distinct workload families.
@@ II-B PauliNet and FermiNet via the deepqmc codebase
Figure 3: Overview of the execution pipeline of Psiformer and Orbformer in the oneqmc codebase. The upper panels show the Orbformer-specific modules used in Stage B, namely the orbital generator, nuclei message-passing neural network (MPNN), and electron Transformer, while the lower block diagram contrasts the Orbformer path with the simpler Psiformer path. Specifically, Stage B (Transformer-based electron representation learning) and Stage C (orbital/readout projection) are the main departures from deepqmc; Stages A and D (feature construction and Slater-determinant wavefunction assembly) follow the same high-level structure as in Figure 2, while Stage E uses a Hutchinson-style Laplacian estimator and is omitted here for brevity.
@@ II-C Psiformer and Orbformer via the oneqmc codebase
  • Stage E (derivative/Laplacian evaluation). Local-energy evaluation includes a kinetic-energy (KE) term $T_{\mathrm{loc}}(\{r\})$ and a potential-energy (PE) term $V(\{r\})$, as shown in Eq. (4). Compared with deepqmc, oneqmc uses a Hutchinson-style Laplacian estimator for the KE term, replacing explicit second-order Laplacian evaluation with one first-order differentiation plus random projection; the PE term is evaluated directly from the sampled coordinates. This keeps the number of differentiation passes constant rather than scaling with all Cartesian directions of all electrons [2].
Summary. For Psiformer/Orbformer in oneqmc, Stages A–D provide forward wavefunction evaluation and Stage E provides local-energy evaluation through a Hutchinson-style Laplacian path, again within the MCMC loop. This execution pattern supports the profiling analysis in Section III.
@@ III-A Profiling Settings
  • Workload configuration. We use 32-bit floating point (FP32) and a batch of 1024 electron configurations (parallel MCMC walkers) for all experiments in this section. For kernel-level profiling, we focus on methane (CH4; 10 electrons). For PauliNet and FermiNet, each profiling run uses default deepqmc sampling settings and consists of ten Metropolis–Hastings MCMC transitions [22, 7], one forward pass through Stages A–D, and $3N_{\mathrm{e}}$ JVP replays for Laplacian evaluation (Stage E), which re-execute computations from Stages A–D. For Psiformer and Orbformer, we use corresponding oneqmc defaults with the same precision and batch settings; each run contains one forward pass through Stages A–D and one Stage E local-energy path based on a Hutchinson-style Laplacian estimator, i.e., one first-order differentiation plus random projection [2], within the same Metropolis–Hastings process.
Unless otherwise stated, all analyses in Sections III-B to III-F use the profiling setup defined in this subsection.
@@ IV Discussion
Section III suggests two coupled facts about NNVMC execution. First, many dominant kernels are low-arithmetic-intensity elementwise/layout operators concentrated in the memory-bound roofline region [36] (Figures 5 and 6), so data movement and kernel granularity can be major bottlenecks. Second, memory-bound dominance does not mean GEMM is negligible: GEMM in embedding/message-passing stages (Stages A–C) still contributes a substantial runtime share and becomes more prominent in higher-capacity models (e.g., FermiNet vs. Psiformer). This combination indicates that NNVMC can behave as a mixed and phase-dependent workload rather than a single-kernel regime. In particular, Stage E derivative/Laplacian evaluation replays earlier computations and amplifies fine-grained kernels, which can shift the effective operator balance across stages and ansätze [2, 29, 6]. Therefore, compute-only or memory-only tuning may be insufficient.
Based on these observations, Section IV discusses five algorithm–hardware co-design directions: PIM for memory-bound kernel clusters, collaborative GPU–PIM partitioning, phase-aligned reconfigurable acceleration, architecture support beyond attention kernels, and CPU/SSD memory offloading [23, 35, 14, 8, 32]. Table I summarizes how Section III findings map to these directions.
@@ IV-D Architecture Support Beyond Attention
In Orbformer, FlashAttention-related kernels account for only 10% to 20% of total end-to-end runtime (Figure 5(d)). Therefore, optimizing attention alone may be unlikely to deliver large end-to-end gains. In our tests, Orbformer with FlashAttention-1 achieves less than 20% speedup over the original attention implementation for one MCMC sampling step. FlashAttention-3 [31], with optimizations such as warp specialization and asynchronous warp-group pipelining, already reports high hardware utilization on modern GPUs. This suggests that additional attention-only tuning may provide diminishing end-to-end returns for NNVMC unless other dominant kernel families are also addressed.
@@ IV-E CPU/SSD Memory Offload
V Related Works
V-A Profiling Workload in Other Emerging AI Applications

### 22 Mar 2026 v2

@@ I Introduction
NNVMC represents the electronic wavefunction with a neural network and optimizes model parameters under the variational principle. Prior studies show that this approach can achieve strong accuracy while offering favorable asymptotic scaling in representative settings (often cited as $O(N^{4})$ versus $O(N^{7})$ for CCSD(T)) [4]. Since Neural Quantum States were introduced, multiple neural wavefunction ansätze have been proposed. In this work, we focus on four representative models for ground-state problems: FermiNet [32], PauliNet [11], Psiformer [39], and Orbformer [6]. Other models, e.g., DeepSolid [25], LapNet [21], QiankunNet [37, 14], Neural Pfaffians [8, 7], and FiRE [35] are out of scope for this survey; in this study, we treat them as extensions of these parent families rather than distinct workload families.
@@ II-A Variational quantum Monte Carlo with neural-network ansätze (NNVMC)
Beyond the released oneqmc models profiled here, other potentially more efficient model designs may also exist. For example, approximating the Laplacian through Hutchinson trace estimation (HTE) is a potential direction for improving the efficiency of local-energy evaluation [13]. However, since no corresponding public model release is currently available, we leave profiling of such designs to future work.
@@ II-B PauliNet and FermiNet via the deepqmc codebase
Figure 3: Overview of the execution pipeline of Psiformer and Orbformer in the oneqmc codebase. The upper panels show the Orbformer-specific modules used in Stage B, namely the orbital generator, nuclei message-passing neural network (MPNN), and electron Transformer, while the lower block diagram contrasts the Orbformer path with the simpler Psiformer path. Specifically, Stage B (Transformer-based electron representation learning) and Stage C (orbital/readout projection) are the main departures from deepqmc; Stages A and D (feature construction and Slater-determinant wavefunction assembly) follow the same high-level structure as in Figure 2, while Stage E uses a forward Laplacian estimator and is omitted here for brevity.
@@ II-C Psiformer and Orbformer via the oneqmc codebase
  • Stage E (derivative/Laplacian evaluation). Local-energy evaluation includes a kinetic-energy (KE) term $T_{\mathrm{loc}}(\{r\})$ and a potential-energy (PE) term $V(\{r\})$, as shown in Eq. (4). Compared with deepqmc, oneqmc uses a forward Laplacian [21] estimator for the KE term, improving the Laplacian computation by directly propagating first-order derivatives and Laplacian terms through the network in a single forward pass without any explicit Hessian constructions; the PE term is evaluated directly from the sampled coordinates. This keeps the number of differentiation passes constant rather than scaling with all Cartesian directions of all electrons [2].
Summary. For Psiformer/Orbformer in oneqmc, Stages A–D provide forward wavefunction evaluation and Stage E provides local-energy evaluation through a forward Laplacian path, again within the MCMC loop. This execution pattern supports the profiling analysis in Section III.
@@ III-A Profiling Settings
  • Workload configuration. We use 32-bit floating point (FP32) and a batch of 1024 electron configurations (parallel MCMC walkers) for all experiments in this section. For kernel-level profiling, we focus on methane (CH4; 10 electrons). For PauliNet and FermiNet, each profiling run uses default deepqmc sampling settings and consists of ten Metropolis–Hastings MCMC transitions [27, 9], one forward pass through Stages A–D, and $3N_{\mathrm{e}}$ JVP replays for Laplacian evaluation (Stage E), which re-execute computations from Stages A–D. For Psiformer and Orbformer, we use corresponding oneqmc defaults with the same precision and batch settings; each run contains one forward pass through Stages A–D and one Stage E local-energy path based on a forward Laplacian estimator [21], within the same Metropolis–Hastings process.
Unless otherwise stated, all analyses in Sections III-CIII-F use the profiling setup defined in this subsection.
@@ IV Discussion
Section III suggests two coupled facts about NNVMC execution. First, many dominant kernels are low-arithmetic-intensity elementwise/layout operators concentrated in the memory-bound roofline region [42] (Figures 5 and 6), so data movement and kernel granularity can be major bottlenecks. Second, memory-bound dominance does not mean GEMM is negligible: GEMM in embedding/message-passing stages (Stages A–C) still contributes a substantial runtime share and becomes more prominent in some higher-capacity model variants (e.g., FermiNet vs. PauliNet). This combination indicates that NNVMC can behave as a mixed and phase-dependent workload rather than a single-kernel regime. In particular, Stage E derivative/Laplacian evaluation replays earlier computations and amplifies fine-grained kernels, which can shift the effective operator balance across stages and ansätze [2, 34, 6]. Therefore, compute-only or memory-only tuning may be insufficient.
Based on these observations, this section discusses five algorithm–hardware co-design directions: PIM for memory-bound kernel clusters, collaborative GPU–PIM partitioning, phase-aligned reconfigurable acceleration, architecture support beyond attention kernels, and CPU/SSD memory offloading [28, 41, 18, 10, 38]. Table I summarizes how findings in Section III map to these directions.
@@ IV-D Architecture Support Beyond Attention
In Orbformer, FlashAttention-related kernels account for a limited share of total runtime (Figure 5(d)). Therefore, optimizing attention alone may be unlikely to deliver large end-to-end gains. In our tests, Orbformer with FlashAttention-1 achieves less than 20% speedup over the original attention implementation for one MCMC sampling step. FlashAttention-3 [36], with optimizations such as warp specialization and asynchronous warp-group pipelining, already reports high hardware utilization on modern GPUs. This suggests that additional attention-only tuning may provide diminishing end-to-end returns for NNVMC unless other dominant kernel families are also addressed.
@@ IV-E CPU/SSD Memory Offload
V Related Work
V-A Profiling Workloads in Other Emerging AI Applications